# Attention-Driven Multi-Agent Reinforcement Learning: Enhancing Decisions with Expertise-Informed Tasks

### v1

@@ Introduction
MPE - Go Landmark
The primary objective is to get closer to a landmark.
Use a PID controller to convert distance and angle from the landmark into discrete actions.
[1, 0, distance, sin(radial), cos(radial), own_velocity]
MPE - Avoid Collision
Collisions result in penalties and propel you to an unexpected position.
Select the closest action to the perpendicular direction of the ally’s direction.
[0, 1, ally_distance, sin(ally_radial), cos(ally_radial), own_velocity]
SISL - Explore
In specific situations, it is better to move in a selected direction to find better conditions.
Direct conversion, as base actions are already directions.
[1, 0, 0, unitary_distance, sin(direction), cos(direction), own_agent_velocity]
SISL - Pursuer Evader
The objective is to capture evaders, necessitating movement towards them. Additionally, it involves splitting up for effective capture.
Select action that minimizes distance to the evader and facilitates splitting from allies when close.
[0, 1, 0, evader_distance, sin(evader_radial), cos(evader_radial), own_agent_velocity]
SISL - Follow Ally
Working alone is not feasible, and allies may have a better understanding of the situation.
Select action that minimizes distance to the allies.
[0, 0, 1, ally_distance, sin(ally_radial), cos(ally_radial), own_velocity]
@@
To validate the proposed methodology, we select the MPE and SISL, two established scenarios focusing on learning collaborative behavior. PettingZoo (?), version 1.24.0, serves as the environment testbed. The MPE Simple Spread is a scenario where agents must efficiently occupy landmarks while avoiding collisions, testing cooperative and spatial-awareness skills. The SISL Pursuit is a predator-prey game in a discrete setting, where agents work together to trap the prey with limited views and random prey moves. We develop extended versions of the base environments, including the Task Generator and the Task to Action Converter, to process observations, generate tasks, and produce valid actions based on the policy selection. Table 1 detail the tasks for the Single Spread and Pursuit environments, respectively, including embedded knowledge, action conversion, and selected features.
@@
To showcase the benefits of our approach, we conduct experiments to define desired performances, comparable to published results. We also establish a random action-based baseline to ensure minimal learning effectiveness and serve as a reference for comparing different environment versions. Baselines are calculated from 30 repetitions of 50 episodes, choosing random uniformly distributed actions at each step.
@@
The evaluation of the Single Spread scenario to establish a benchmark yields varied results. Among the algorithms, four (IPPO, IQL, VDN, and ISAC) demonstrate comparable performances, while three (MAPPO, MASAC and QMIX) fail to find effective policies. This underperformance can be attributed to the use of default parameters without fine-tuning. The results of the top performing four algorithms, however, align well with leading results reported in the literature, validating their use as a reliable benchmark for this study. The highest mean reward is recorded by IPPO at -8.67 (CI: -9.48, -7.90) in after iteration 425, which is even the best-known result for the adopted configuration. We also calculate the mean reward for the tests in the last 50 steps, with the best result for IPPO being -9.81 (CI: -9.94, -9.67). Regarding the SISL Pursuit, the benchmark achieves a best reward of 633.6 (CI: 597.2, 668.3) which aligns with the best results found in the literature. We also watch simulations conducted by the best policy in both cases and confirm visually that the agents collaborate in a very efficient manner.
@@
This study introduces an alternative approach to MARL, integrating domain knowledge through higher-level tasks and attention-based policy mechanisms. Experiments across standard MARL scenarios demonstrate our methodology’s potential, surpassing benchmarks in learning efficiency and behavioral effectiveness. The inherent flexibility of our approach enables scalable adaptation to more complex scenarios and variable input observations. It preserves cooperative capabilities without additional training. Our method relies on the careful design of higher-level tasks, which can either limit or bolster the model’s learning potential. Personalizing these tasks to specific problems underscores the need for customization in each application which is not always a feasible approach. When customization is a feasible, however, it can significantly simplify the overall learning process. Future work will explore the benefits of our methodology in more realistic settings, such as autonomous vehicle systems and military operations, to further assess its potential in challenging real-world applications.
@@ Acknowledgments
This work was supported by the National Council for Scientific and Technological Development—CNPq 307691/2020-9; the Research and Projects Financing (FINEP—The Brazilian Innovation Agency) 2824/20; and the Brazilian Air Force Postgraduate Program in Operational Applications (PPGAO).

### v3

@@ Introduction
MPE - Go Landmark The primary objective is to get closer to a landmark. Use a PID controller to convert distance and angle from the landmark into discrete actions. [1, 0, distance, sin(radial), cos(radial), own_velocity]
MPE - Avoid Collision Collisions result in penalties and propel you to an unexpected position. Select the closest action to the perpendicular direction of the ally’s direction. [0, 1, ally_distance, sin(ally_radial), cos(ally_radial), own_velocity]
SISL - Explore In specific situations, it is better to move in a selected direction to find better conditions. Direct conversion, as base actions are already directions. [1, 0, 0, unitary_distance, sin(direction), cos(direction), own_agent_velocity]
SISL - Pursuer Evader The objective is to capture evaders, necessitating movement towards them. Additionally, it involves splitting up for effective capture. Select action that minimizes distance to the evader and facilitates splitting from allies when close. [0, 1, 0, evader_distance, sin(evader_radial), cos(evader_radial), own_agent_velocity]
SISL - Follow Ally Working alone is not feasible, and allies may have a better understanding of the situation. Select action that minimizes distance to the allies. [0, 0, 1, ally_distance, sin(ally_radial), cos(ally_radial), own_velocity]
@@
To validate the proposed methodology, we select the MPE and SISL, two established scenarios focusing on learning collaborative behavior. PettingZoo (?), version 1.24.0, serves as the environment testbed. The MPE Simple Spread is a scenario where agents must efficiently occupy landmarks while avoiding collisions, testing cooperative and spatial-awareness skills. The SISL Pursuit is a predator-prey game in a discrete setting, where agents work together to trap the prey with limited views and random prey moves. We develop extended versions of the base environments, including the Task Generator and the Task to Action Converter, to process observations, generate tasks, and produce valid actions the policy selection. Table 1 detail the tasks for the Single Spread and Pursuit, including embedded knowledge, action conversion, and selected features.
@@
To showcase the benefits of our approach, we conduct experiments to define desired performances, comparable to published results. We also establish a random action-based baseline to ensure minimal learning effectiveness and serve as a reference for comparing different environments. Baselines are calculated from 30 repetitions of 50 episodes, choosing random uniformly distributed actions at each step.
@@
The evaluation of the Single Spread scenario to establish a benchmark yields varied results. Among the algorithms, four (IPPO, IQL, VDN, and ISAC) demonstrate comparable performances, while three (MAPPO, MASAC and QMIX) fail to find effective policies. This underperformance can be attributed to the use of default parameters without fine-tuning. The results of the top performing four algorithms, however, align well with leading results reported in the literature, validating their use as a reliable benchmark. The highest mean reward is recorded by IPPO at -8.67 (CI: -9.48, -7.90) in iteration 425, which is even the best-known result for the adopted configuration. We also calculate the mean reward for the tests in the last 50 steps, with the best result for IPPO being -9.81 (CI: -9.94, -9.67). Regarding the SISL Pursuit, the benchmark achieves a best reward of 633.6 (CI: 597.2, 668.3) which aligns with the best results in the literature. We also watch simulations of the best policies and confirm visually that the agents collaborate in a very efficient manner.
@@
This study introduces an alternative MARL approach, integrating domain knowledge through higher-level tasks and attention-based policy. Experiments across standard MARL scenarios demonstrate our methodology’s potential, surpassing benchmarks in learning efficiency and effectiveness. The inherent flexibility enables scalable adaptation to more complex scenarios and variable observations while preserving cooperative capabilities without additional training. Our method relies on designing higher-level tasks, which can either limit or bolster the model’s potential. Personalizing tasks to specific problems underscores the need for customization in each application, which is not always feasible. When customization is feasible, however, it can significantly simplify the overall learning process. Future work will explore our methodology’s benefits in more realistic settings, such as autonomous vehicles and military operations, to further assess its potential in real-world applications.
@@ Acknowledgments
This work was supported by the National Council for Scientific and Technological Development—CNPq 307691/2020-9 and the Brazilian Air Force Postgraduate Program in Operational Applications (PPGAO).